OpenStudy (anonymous):

Identify the type of reasoning in the following situation: You notice a pattern in the following sequence of numbers: -1, 5, 11, 17, 23… You conclude that the next number must be 29.

OpenStudy (anonymous):

+6 ?

OpenStudy (anonymous):

You're always adding 6 :P

OpenStudy (anonymous):

-1 +6 = 5 5 + 6 = 11 11 + 6 = 17 17 + 6 = 23 23 + 6 = 29 :P

OpenStudy (anonymous):

i mean the type of reasoning lol